OpenStudy (code45):

Given the frequencies listed below, calculate the corresponding periods. a. 24 Hz b. 8 MHz c. 140 KHz

OpenStudy (anonymous):

Just take the resprocal of the given number.

ganeshie8 (ganeshie8):

hint : \(period = \frac{1}{frequency}\)

OpenStudy (code45):

period=1/f so 1/24Hz=.0417

ganeshie8 (ganeshie8):

.0417 seconds = 41.7 milli seconds

ganeshie8 (ganeshie8):

dont let MHz and KHz confuse you. MHz just means, Mega Hertz 1 MHz = 10^6 Hz

OpenStudy (code45):

2nd question 1/8MHz = 0.000000125=.125*10^-6s=0.125us

ganeshie8 (ganeshie8):

Perfect ! same thing you can write like this aswell : 0.125 us = 125 ns

ganeshie8 (ganeshie8):

np... yw ! :)

OpenStudy (code45):

c. 140KHz t=1/f =1/140KHz = 0.00000714s = 7.14*10^-6s =7.14us

ganeshie8 (ganeshie8):

correct :D 140 KHz = .140 MHz = 1/(.14 * 10^6) s = 7.14 us